4x/5+3/2=2x
How do you work out this problem?
I just can't seem to get it.
1/6-x/2=x-5/3
I don't understand this one either...please help me.

Answer by stanbon(75887)

4x/5+3/2=2x
How do you work out this problem?
I just can't seem to get it.
1/6-x/2=x-5/3
I don't understand this one either...please help me.
========================================================
4x/5+3/2=2x
---
Multiply thru by 10 to get:
8x + 15 = 20x
12x = 15
x = 5/4
================
1/6-x/2=x-5/3
Multiply thru by 6 to get:
1 - 3x = 6x - 10
9x = 11
x = 11/9
================
Cheers,
Stan H.